OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is odd or 1 die shows a 4? A 6 X 6 table of dice outcomes will help you to answer this question. Write fractions using the slash ( / ) key. Reduce fractions to their lowest terms.

OpenStudy (anonymous):

Total number of possible outcomes = 6^2 = 36 Sum of the digits---No. of outcomes 02 ------------------------- 01 03 ------------------------- 02 04 ------------------------- 03 05 ------------------------- 04 06 ------------------------- 05 07 ------------------------- 06 08 ------------------------- 05 09 ------------------------- 04 10 ------------------------- 03 11 ------------------------- 02 12 ------------------------- 01 Total --------------------- 36 P(sum of the digits is odd) = P(A) = (2+4+6+4+2) / 36 = 18/36 P(sum of the digits > 6) = P(B) = (6+5+4+3+2+1) / 36 = 21/36 P(A and B) = (6+4+2) / 36 = 12/36 Required probability = P(A or B) = P(A)+P(B)-P(A and B) =18/36 + 21/36 - 12/36 = 27/36 = 3/4

OpenStudy (anonymous):

now did that help?

OpenStudy (anonymous):

so complicated but yes, i think so :)

OpenStudy (anonymous):

well u askd Write fractions using the slash ( / ) key. Reduce fractions to their lowest terms. right?

OpenStudy (anonymous):

so u sumed up the numbers as in lowest terms"